On Earth, the gravitational field intensity is 9.8 N/kg. Mercury has a mass of 3.28 x 1023 kg and a radius of 2.57 x 106 m. Calculate the gravitational field intensity on Mercury in N/kg.




1
Expert's answer
2021-07-12T12:16:39-0400
"g=\frac{GM}{R^2}\\g=\frac{(6.67\cdot10^{-11})(3.28\cdot10^{23})}{(2.57\cdot10^{6})^2}=3.3\frac{N}{kg}"